# Efficient calculation of magnetic fields from ferromagnetic materials near strong electromagnets, and application to stellarator coil optimization

### v1

@@ 4.3 Example
To illustrate the method with an example, we return to the quasi-isodynamic stellarator and steel geometry shown in figure 3, and explained in section 3. The magnetic field for the initial state is displayed in figure 7. The eight panels all show aspects of the magnetic field on the target plasma surface, i.e. the surface that was used in the initial optimization of the coil shapes in which the ferromagnetic components were neglected. The plasma configuration has a ratio of plasma pressure to magnetic pressure of $\beta=1.7\%$, so the field due to the plasma current is non-negligible compared to coil ripple on the plasma surface, and is computed using the virtual casing method [malhotra2019efficient]. However, throughout the steel region, the ratio $|\mathbf{B}^{plasma}|/|\mathbf{B}^{coils}|$ is found to be quite small, $2.4\%$, so $|\mathbf{B}^{plasma}|$ is neglected for determining the steel magnetization in the rest of this paper. Indeed, the effect of plasma current on the steel field would be a small correction to an already small correction. The four columns show the total field and the separate contributions from coils, plasma current, and the steel. The top row shows the magnitude of the field in Tesla, while the bottom row shows the field component normal to the target plasma surface, as this normal component highlights the errors in producing the desired flux surface. It can be seen that $\mathbf{B}^{ferr}$ is quite small, less than 1% of the total field. This fact supports our approximation that $\mathbf{B}^{ferr}$ can be neglected when computing the direction of the magnetization, and suggests that the effect of $\mathbf{B}^{ferr}$ on the plasma properties is unlikely to be dramatic. Nonetheless, looking at the components normal to the target plasma surface in the bottom row, $\mathbf{B}^{ferr}\cdot\mathbf{n}$ is comparable to the modular coil ripple and $\mathbf{B}^{plasma}\cdot\mathbf{n}$. Therefore $\mathbf{B}^{ferr}$ should not be ignored completely, and we can expect it may be large enough to have some effect on certain physics properties of the plasma. As noted above, the largest magnetic perturbation from the steel is in the region under the ports.
@@
Free-boundary MHD equilibria are computed for these initial coils, both without and with the steel. The equilibria are computed using the code VMEC [hirshman1986three]. For the free-boundary calculation with $\mathbf{B}^{ferr}$, the field from the dipoles representing the steel is evaluated on a tensor-product grid in cylindrical coordinates and saved in VMEC’s standard MGRID input format.
Figure 8 shows the profiles of rotational transform $\iota$ for the original fixed-boundary target configuration as well as for free-boundary equilibria both excluding and including the field from the steel. If the steel field is excluded, the coils do a good job matching the $\iota$ profile of the target configuration. However when the steel field is included, there is a systematic decrease in the $\iota$ profile. This $\iota$ downshift due to ferromagnetic material has been noted previously [harmeyer1999effect, ji2017investigation]. It can be understood from analysis of a straight cylinder, where the ferromagnetic layers reduce the poloidal field while not affecting the toroidal field [harmeyer1999effect].
For stellarator designs with an island divertor, such as W7-X and the configuration here (which uses the $\iota=3/4$ resonance just outside the plasma), this shift to $\iota$ will affect the locations of heat strike points on the divertor, so it must be accounted for. This effect is seen in figure 9, which show Poincare plots at the operating $\beta=1.7\%$ computed using the code HINT [suzuki2006development, suzuki2017hint]. The Poincare plots at $\beta=0$ ($\mathbf{B}^{coils}+\mathbf{B}^{ferr}$ only) are nearly identical. The original coils were designed to produce a chain of four islands at the plasma edge (top panel). When the field from the steel is included (middle panel), the islands are shifted outward from the domain. Also a small higher-harmonic island has arisen due to the field perturbation. While both of these issues could be corrected with additional control coils, we can address them directly with small adjustments to the modular coil shapes.
@@
Here, $\int_{\partial P}d^{2}a$ is a surface integral over the plasma boundary, $N$ is the number of unique coil shapes, $j$ is an index over these shapes, $L_{j}$ is the length of coil $j$, $\int_{j}d\ell$ is an integral over coil $j$ with respect to arclength, $\kappa(\ell)$ is the curvature, $\mathbf{r}_{j}(\ell)$ is the position vector along coil $j$, and $\mathbf{r}_{S}$ is a position vector on this surface. The quantity $b^{res}_{k}$ is the $k$-th harmonic resonant field, eq (14) in [boozer2004physics], associated with the edge island chain. The quantities $L_{*}$, $\kappa_{*}$, $d_{cc*}$, $d_{cs*}$, and $b^{res}_{k*}$ are constant target values, and are set to the same values used to generate the original coils in figure 3. Therefore the coil lengths, maximum curvature, minimum coil-coil distance, and minimum coil-plasma distance are practically unchanged during the optimization that includes the steel. Other than the weights $w_{k}$ on the resonances, no additional weights in (4.3) are required. The coils before and after the optimization are compared in figure 10. It can be seen that the changes to the coil shapes are relatively minor, which is not surprising given that the field from the steel is small compared to the total field. The most noticeable changes to the coil shapes are in the regions on the outboard midplane near the ports.
@@
Figure 11 shows the contributions to $\mathbf{B}$, and their components normal to the target plasma boundary, following the optimization. Comparing figures 7 and 11, the right columns show that $\mathbf{B}^{ferr}$ is nearly unchanged, reflecting the fact that the change to the coil shapes is small. Looking at the sixth panel ($\mathbf{B}^{coils}\cdot\mathbf{n})/|\mathbf{B}^{total}|$, focusing on the lower left corner, it can be seen that the optimized coils generate an approximately equal and opposite ripple to the ripple from the steel near the ports. As a result, comparing the lower-left panels of figures 7 and 11, the total ripple near the ports is significantly reduced.
@@
We also find that the field from the ferritic material has negligible effect on several measures of confinement. The neoclassical transport measure $\epsilon_{eff}$ is shown in figure 14 for the three free-boundary configurations. No systematic difference is apparent. Figure 15 shows the results of guiding-center trajectory calculations in the three free-boundary configurations. For these calculations, conducted using the code SIMPLE [albert2020accelerated, albert2020symplectic], 5000 particles are initialized on a flux surface, and followed without collisions for 0.1 seconds. The original size and field strength of the configurations are used for the tracing. In all three configurations, no particles are lost if the initial surface $s$ (normalized toroidal flux) is <0.5<0.5, and losses from $s=0.5$ are $0.9\%$. For particles initialized farther out (ss= 0.6-0.7), losses may be slightly higher for the original coils with steel, but the effect is small. Overall, we conclude that the field from the steel components does not significantly affect these two neoclassical confinement measures.
@@
Figure 15: The steel is found to have only a minor effect on confinement of guiding-center trajectories.
@@ Appendix A Force on a magnetized region
Forces on magnetized objects can be computed in a wide variety of ways: integration of the Maxwell stress over a surface, differentiating an energy (giving the so-called Korteweg–Helmholtz force), equivalent magnetic charges, or the Lorentz force on magnetization currents or dipoles [Melcher, deMedeiros1999distribution, engel2002electromagnetic, jackson, bobbio2000equivalent]. Some of these methods can be expressed either in terms of the total magnetic field or the part from sources outside the magnetized domain [bobbio2000equivalent]. Here we demonstrate the equivalence of two intuitive methods for computing the force, the magnetization current and dipole methods. Let $\mathbf{f}$ denote the total force on a magnetized region $\Omega$. This force can be understood to originate from the Lorentz force on the magnetization currents:
@@
where $\mathbf{J}_{b}=\nabla\times\mathbf{\mathbf{M}}$ is the current density associated with bound currents and $\mathbf{K}_{b}=\mathbf{M}\times\mathbf{n}$ is the surface current density associated with bound currents, with $\mathbf{n}$ the outward unit normal of $\Omega$. In (11), $\mathbf{B}$ can be taken to be $\mathbf{B}^{ext}$, the field from only sources outside $\Omega$, namely $\mathbf{B}^{coils}+\mathbf{B}^{plasma}$, since the field from the magnetized material cannot exert any net force on itself [jackson, landau]. To simplify notation the ext superscript is omitted in this appendix. We apply $(\nabla\times\mathbf{M})\times\mathbf{B}=-\nabla(\mathbf{M}\cdot\mathbf{B})+(\nabla\mathbf{B})\cdot\mathbf{M}+\mathbf{B}\cdot\nabla\mathbf{M}$ to the first term of (11), and apply the BAC-CAB rule to the second term. Using $\int_{\Omega}d^{3}r\nabla g=\int_{\partial\Omega}d^{2}r\,\mathbf{n}g$ for $g=\mathbf{M}\cdot\mathbf{B}$, a cancellation occurs to leave

### v2

@@ 4.3 Example
To illustrate the method with an example, we return to the quasi-isodynamic stellarator and steel geometry shown in figure 3, and explained in section 3. The magnetic field for the initial state is displayed in figure 8. The eight panels all show aspects of the magnetic field on the target plasma surface, i.e. the surface that was used in the initial optimization of the coil shapes in which the ferromagnetic components were neglected. The plasma configuration has a ratio of plasma pressure to magnetic pressure of $\beta=1.7\%$, so the field due to the plasma current is non-negligible compared to coil ripple on the plasma surface, and is computed using the virtual casing method [32]. However, throughout the steel region, the ratio $|\mathbf{B}^{plasma}|/|\mathbf{B}^{coils}|$ is found to be quite small, $2.4\%$, so $|\mathbf{B}^{plasma}|$ is neglected for determining the steel magnetization in the rest of this paper. Indeed, the effect of plasma current on the steel field would be a small correction to an already small correction. The four columns show the total field and the separate contributions from coils, plasma current, and the steel. The top row shows the magnitude of the field in Tesla, while the bottom row shows the field component normal to the target plasma surface, as this normal component highlights the errors in producing the desired flux surface. It can be seen that $\mathbf{B}^{ferr}$ is quite small, less than 1% of the total field. This fact supports our approximation that $\mathbf{B}^{ferr}$ can be neglected when computing the direction of the magnetization, and suggests that the effect of $\mathbf{B}^{ferr}$ on the plasma properties is unlikely to be dramatic. Nonetheless, looking at the components normal to the target plasma surface in the bottom row, $\mathbf{B}^{ferr}\cdot\mathbf{n}$ is comparable to the modular coil ripple and $\mathbf{B}^{plasma}\cdot\mathbf{n}$. Therefore $\mathbf{B}^{ferr}$ should not be ignored completely, and we can expect it may be large enough to have some effect on certain physics properties of the plasma. As noted above, the largest magnetic perturbation from the steel is in the region corresponding to the ports.
@@
Free-boundary MHD equilibria are computed for these initial coils, both without and with the steel. The equilibria are computed using the code VMEC [21]. For the free-boundary calculation with $\mathbf{B}^{ferr}$, the field from the dipoles representing the steel is evaluated on a tensor-product grid in cylindrical coordinates and saved in VMEC’s standard MGRID input format.
Figure 9 shows the profiles of rotational transform $\iota$ for the original fixed-boundary target configuration as well as for free-boundary equilibria both excluding and including the field from the steel. If the steel field is excluded, the coils do a good job matching the $\iota$ profile of the target configuration. However when the steel field is included, there is a systematic decrease in the $\iota$ profile. This $\iota$ downshift due to ferromagnetic material has been noted previously [17, 24]. It can be understood from analysis of a straight cylinder, where the ferromagnetic layers reduce the poloidal field while not affecting the toroidal field [17].
For stellarator designs with an island divertor, such as W7-X and the configuration here (which uses the $\iota=3/4$ resonance just outside the plasma), this shift to $\iota$ will affect the locations of heat strike points on the divertor, so it must be accounted for. This effect is seen in figure 10, which show Poincare plots at the operating $\beta=1.7\%$ computed using the code HINT [40, 41]. The Poincare plots at $\beta=0$ ($\mathbf{B}^{coils}+\mathbf{B}^{ferr}$ only) are nearly identical. The original coils were designed to produce a chain of four islands at the plasma edge (top panel). When the field from the steel is included (middle panel), the islands are shifted outward from the domain. Also a small higher-harmonic island has arisen due to the field perturbation. While both of these issues could be corrected with additional control coils, we can address them directly with small adjustments to the modular coil shapes.
@@
Here, $\int_{\partial P}d^{2}a$ is a surface integral over the plasma boundary, $N$ is the number of unique coil shapes, $j$ is an index over these shapes, $L_{j}$ is the length of coil $j$, $\int_{j}d\ell$ is an integral over coil $j$ with respect to arclength, $\kappa(\ell)$ is the curvature, $\mathbf{r}_{j}(\ell)$ is the position vector along coil $j$, and $\mathbf{r}_{S}$ is a position vector on this surface. The quantity $b^{res}_{k}$ is the $k$-th harmonic resonant field, eq (14) in [5], associated with the edge island chain:
$b_{k}^{res}=\frac{1}{2\pi^{2}}\int d\vartheta\int d\phi\frac{\mathbf{B}\cdot\nabla\psi}{\mathbf{B}\cdot\nabla\phi}\sin(k[4\vartheta-3\phi]),$} (11)
for toroidal flux $\psi$ and straight-field-line poloidal angle $\vartheta$. In (4.3), the second and third harmonics ($k=2,3$) are controlled in addition to the fundamental ($k=1$) to prevent the appearance of higher-order island chains. The quantities $L_{*}$, $\kappa_{*}$, $d_{cc*}$, $d_{cs*}$, and $b^{res}_{k*}$ are constant target values, and are set to the same values used to generate the original coils in figure 3. Therefore the coil lengths, maximum curvature, minimum coil-coil distance, and minimum coil-plasma distance are practically unchanged during the optimization that includes the steel. Other than the weights $w_{k}$ on the resonances, no additional weights in (4.3) are required. The coils before and after the optimization are compared in figure 11. It can be seen that the changes to the coil shapes are relatively minor, which is not surprising given that the field from the steel is small compared to the total field. The most noticeable changes to the coil shapes are in the regions on the outboard midplane near the ports.
@@
Figure 12 shows the contributions to $\mathbf{B}$, and their components normal to the target plasma boundary, following the optimization. Comparing figures 8 and 12, the right columns show that $\mathbf{B}^{ferr}$ is nearly unchanged, reflecting the fact that the change to the coil shapes is small. Looking at the sixth panel ($\mathbf{B}^{coils}\cdot\mathbf{n})/|\mathbf{B}^{total}|$, focusing on the lower left corner, it can be seen that the optimized coils generate an approximately equal and opposite ripple to the ripple from the steel near the ports. As a result, comparing the lower-left panels of figures 8 and 12, the total ripple near the ports is significantly reduced.
@@
We also find that the field from the ferritic material has negligible effect on several measures of confinement. The neoclassical transport measure $\epsilon_{eff}$ is shown in figure 15 for the three free-boundary configurations. No systematic difference is apparent. Next, figure 16 shows the results of guiding-center trajectory calculations in the three free-boundary configurations. For these calculations, conducted using the code SIMPLE [1, 2], first 25,000 particles are initialized throughout the plasma volume proportional to the fusion reaction rate from the local density and temperature, and followed without collisions for 0.1 seconds. The original size and field strength of the configurations are used for the tracing. In all three configurations, losses are extremely low, $0.3\%$, without any significant effect from the steel. It is conceivable that differences might be greater if particles are initialized at larger minor radius, closer to the steel. To check this hypothesis, further calculations are done with particles initialized on specific flux surfaces, with results shown in figure 17. No particles are lost if the initial surface $s$ (normalized toroidal flux) is <0.5<0.5, and losses from $s=0.5$ are $0.9\%$. For particles initialized farther out (ss= 0.6-0.7), losses may be slightly higher for the original coils with steel, but the effect is small. Overall, we conclude that the field from the steel components does not significantly affect either the effective ripple $\epsilon_{eff}$ or energetic particle confinement.
@@
Figure 16: The steel is found to not have a significant effect on confinement of guiding-center trajectories.
Figure 17: Even for guiding center trajectories initialized close enough to the plasma edge to be lost, the steel is found to have only a minor effect on confinement.
@@ Appendix A Force on a magnetized region
Forces on magnetized objects can be computed in a wide variety of ways: integration of the Maxwell stress over a surface, differentiating an energy (giving the so-called Korteweg–Helmholtz force), equivalent magnetic charges, or the Lorentz force on magnetization currents or dipoles [33, 10, 13, 23, 4]. Some of these methods can be expressed either in terms of the total magnetic field or the part from sources outside the magnetized domain [4]. Here we demonstrate the equivalence of two intuitive methods for computing the force, the magnetization current and dipole methods. Let $\mathbf{f}$ denote the total force on a magnetized region $\Omega$. This force can be understood to originate from the Lorentz force on the magnetization currents:
@@
where $\mathbf{J}_{b}=\nabla\times\mathbf{\mathbf{M}}$ is the current density associated with bound currents and $\mathbf{K}_{b}=\mathbf{M}\times\mathbf{n}$ is the surface current density associated with bound currents, with $\mathbf{n}$ the outward unit normal of $\Omega$. In (12), $\mathbf{B}$ can be taken to be $\mathbf{B}^{ext}$, the field from only sources outside $\Omega$, namely $\mathbf{B}^{coils}+\mathbf{B}^{plasma}$, since the field from the magnetized material cannot exert any net force on itself [23, 27]. To simplify notation the ext superscript is omitted in this appendix. We apply $(\nabla\times\mathbf{M})\times\mathbf{B}=-\nabla(\mathbf{M}\cdot\mathbf{B})+(\nabla\mathbf{B})\cdot\mathbf{M}+\mathbf{B}\cdot\nabla\mathbf{M}$ to the first term of (12), and apply the BAC-CAB rule to the second term. Using $\int_{\Omega}d^{3}r\nabla g=\int_{\partial\Omega}d^{2}r\,\mathbf{n}g$ for $g=\mathbf{M}\cdot\mathbf{B}$, a cancellation occurs to leave